# Replica Thermodynamic Trade-off Relations: Entropic Bounds on Network Diffusion and Trajectory Observables

### v1

@@ Results
The bounds given in Eqs. (12) and(14) can be generalized to quantum scenarios using the continuous measurement formalism. Let $\rho(t)$ denote the density operator of the system at time $t$. We assume that its dynamics are governed by a Lindblad master equation:
@@
We now consider continuous measurement in the Lindblad framework, corresponding to continuous monitoring of the environment coupled to the system (see Ref. [57]). In this setting, the system’s stochastic evolution is conditioned on the measurement outcomes registered in the environment, and individual realizations of the dynamics are described by quantum trajectories. The measurement record consists of the types of jump events and their time stamps, which represent the full history of detected quanta (such as emitted photons or transferred particles) over the time interval. Suppose that within the interval $0,\tau]$ there are $J$ jump events, and let $m_{j}\in\{1,\dots,\mathfrak{N}\}$ denote the type of the $j$-th jump occurring at time $t_{j}$. The full measurement record is
$\Gamma_{\mathrm{q}}\equiv\big[(t_{1},m_{1}),(t_{2},m_{2}),\ldots,(t_{J},m_{J})\big],$ (25)
@@
where the subscript $\mathrm{q}$ is used to emphasize that $\Gamma_{\mathrm{q}}$ is a quantum trajectory. The sequence $\Gamma_{\mathrm{q}}$ characterizes one realization of the monitored dynamics and is the basic object from which we construct trajectory-dependent observables. Consider a trajectory observable $N_{\mathrm{q}}(\Gamma_{\mathrm{q}})$. Let $\Gamma_{\mathrm{q}\varnothing}$ be arbitrary quantum trajectory, where there is no jump within the interval $0,\tau]$. Analogous to the condition of Eq. (11), we assume that $N_{\mathrm{q}}(\Gamma_{\mathrm{q}})$ satisfies the following condition:
$N_{\mathrm{q}}(\Gamma_{\mathrm{q}\varnothing})=0.$ (26)
$N_{\mathrm{q}}$ includes counting observables, which simply count the number of jumps. In a similar manner to the classical case, we consider $K$ replicas of the identical quantum Markov process. A quantum trajectory of $K=2$ replicas is
$\Gamma_{\mathrm{q}}^{(\otimes 2)}=\Gamma_{\mathrm{q}}^{(\otimes 2)}\left(\Gamma_{\mathrm{q}}^{(1)},\Gamma_{\mathrm{q}}^{(2)}\right),$ (27)
where $\Gamma_{\mathrm{q}}^{(1)}$ and $\Gamma_{\mathrm{q}}^{(2)}$ trajectories in the first and the second processes, respectively. It is straightforward to generalize to $K$-replica scenario, in which the corresponding trajectory is expressed as $\Gamma_{\mathrm{q}}^{(\otimes K)}$.
It is possible to consider a quantum generalization of the dynamical activity given in Eq. (13). The classical dynamical activity [Eq. (13)] quantifies the extent of activity of the dynamics. For classical Markov processes, the state change occurs only when there are jumps. However, for the quantum scenario, the state change can occur even without jumps due to the coherent contribution arising from $i[H,\rho]$ term in Eq. (23). Let $\mathcal{B}(\tau)$ be the quantum dynamical activity within the interval $0,\tau]$. The quantum dynamical activity plays a central role in trade-off relations in quantum thermodynamics [32, 44, 58, 59, 60]. For the explicit expression of $\mathcal{B}(\tau)$, please see the Methods. It can be shown that the trajectory observable $N_{\mathrm{q}}(\Gamma_{\mathrm{q}})$ satisfies the following relation (see Methods):
@@
$\frac{\mathbb{E}\left[|N_{\mathrm{q}}^{(\otimes K)}|^{s}\right]^{r/(s-r)}}{\mathbb{E}\left[|N_{\mathrm{q}}^{(\otimes K)}|^{r}\right]^{s/(s-r)}}\geq\left(1-\cos\left[\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{\mathcal{B}(t)}}{t}dt\right]^{2K}\right)^{-1},$ (28)
where $0\leq(1/2)\int_{0}^{\tau}\sqrt{\mathcal{B}(t)}/t\,dt\leq\pi/2$. Similarly to the classical counterpart [Eq. (12)], the right-hand side of Eq. (28) depends on the quantum dynamical activity $\mathcal{B}(\tau)$. For the classical case, the right-hand side depends solely on the initial dynamical activity $\mathfrak{a}(t=0)$. On the other hand, $\mathcal{B}(\tau)$ depends on dynamics within the interval $0,\tau]$. Again, by considering $r=1$ and $s\to\infty$ in Eq. (28) we obtain
@@
$\mathbb{E}\left[|N_{\mathrm{q}}^{(\otimes K)}|\right]\leq N_{\mathrm{q}\max}^{(\otimes K)}\left(1-\cos\left[\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{\mathcal{B}(t)}}{t}dt\right]^{2K}\right),$ (29)
where $N_{\mathrm{q}\max}^{(\otimes K)}\equiv\max_{\Gamma_{\mathrm{q}}^{(\otimes K)}}|N_{\mathrm{q}}^{(\otimes K)}(\Gamma_{\mathrm{q}}^{(\otimes K)})|$ and $0\leq(1/2)\int_{0}^{\tau}\sqrt{\mathcal{B}(t)}/t\,dt\leq\pi/2$.
Let us consider the probability distribution of $N_{\mathrm{q}}(\Gamma)$. Similarly to the classical case, we can derive upperbound to the Rényi entropy of $P(N_{\mathrm{q}})$. From Eq. (29), we obtain the following relation (see Methods):
@@
$H_{\alpha}[P(N_{\mathrm{q}})]\leq\frac{2\alpha}{1-\alpha}\ln\cos\left[\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{\mathcal{B}(t)}}{t}dt\right],$ (30)
where $\alpha=2,3,\cdots$. Equation (30) is a quantum generalization of Eq. (17). The left-hand side of Eq. (17) is the Rényi entropy of the observable $N_{\mathrm{q}}$, while the right-hand side depends on the quantum dynamical activity $\mathcal{B}(\tau)$. For integer values of $\alpha$, $\alpha\in\{2,3,\cdots\}$, the relation is exact. Since the inequality of Eq. (17) formally holds for the real value of $\alpha>1$, we conjecture that this relation holds for arbitrary $\alpha>1$.
Figure 4: Numerical verification of the Rényi entropy bound in Eq. (16). The Rényi entropy $H_{\alpha}[p_{\mu}(\tau)]$ versus $\frac{\alpha\tau}{\alpha-1}\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}$ is plotted for (a) $\alpha=1.5$, (b) $\alpha=2$, and (c) $\alpha=3$. Random realizations are shown with points and the dashed line indicates the equality case of Eq. (16). In (a)-(c), the simulations begin by selecting the number of state $D$ between 33 and $50$. Then, the transition rate $W_{\mu^{\prime}\mu}$ is randomly produced, and $\tau$ is chosen within the range of $0.1$ to 1010.
@@
Figure 5: Numerical verification of the Rényi entropy bound in Eq. (17). The Rényi entropy $H_{\alpha}[P(N_{\mathrm{c}})]$ versus $\alpha\tau\mathfrak{a}(t=0)/(\alpha-1)$ is plotted for (a) $\alpha=1.5$, (b) $\alpha=2$, and (c) $\alpha=3$. Random realizations are shown with points and the solid line indicates the equality case of Eq. (17). In (a)-(c), classical simulations begin by selecting the number of states $D$ between 22 and $5$. Then, the transition rate $W_{\mu^{\prime}\mu}$ is randomly produced, and $\tau$ is chosen within the range of $0.1$ to 1010. The observable $N_{\mathrm{c}}(\Gamma)$ is the number of total jumps within $0,\tau]$. The Rényi entropy of the observable is estimated by averaging the results of $10^{3}$ simulations.
Discussion
We have introduced replica Markov processes as a tool to derive thermodynamic trade-off relations for nonlinear functions of probability distributions. Entropic quantities such as Rényi entropies are nonlinear and are therefore harder to handle. By considering $K$ replica processes, we derived a replica trade-off relation that lower-bounds relative moments of trajectory observables by a term expressed by the dynamical activity [Eq. (12)].
We then showed how this replica trade-off relation can be converted into entropic bounds for a single Markov process. First, for a continuous-time random walk on a network, we obtained an upper bound on the Rényi entropy of the position distribution at time $\tau$ [Eq. (16)]. Remarkably, the bound depends only on $\tau$ and the local escape rate from the initial node, i.e., the sum of transition rates to nodes reachable in one hop. This provides an information-theoretic constraint on network diffusion that is computable from purely local data, independent of the global network size or structure. Second, for trajectory observables, we derived an upper bound on the Rényi entropy of the observable distribution [Eq. (17)] in terms of the dynamical activity. This gives an entropic characterization of trajectory-level uncertainty that complements and conceptually generalizes variance-based thermodynamic uncertainty relations.
We also discussed a direct quantum extension within the continuous-measurement quantum trajectory framework. The resulting replica inequality involves the quantum dynamical activity, capturing both jump and coherent contributions, and leads to an entropic bound for quantum counting observables [Eq. (28)].
Methods
.1 Derivation of Eq. (12)

### v2

@@ II.7 Quantum replica Markov process
The bounds given in Eqs. (14) and (17) can be generalized to quantum scenarios using the continuous measurement formalism.
Let $\rho(t)$ denote the density operator of the system at time $t$. We assume that its dynamics are governed by a Lindblad master equation:
@@
We now consider continuous measurement in the Lindblad framework, corresponding to continuous monitoring of the environment coupled to the system (see Ref. [63]). In this setting, the system’s stochastic evolution is conditioned on the measurement outcomes registered in the environment, and individual realizations of the dynamics are described by quantum trajectories. The measurement record consists of the types of jump events and their time stamps, which represent the full history of detected quanta (such as emitted photons or transferred particles) over the time interval. Suppose that within the interval $0,\tau]$ there are $J$ jump events, and let $m_{j}\in\{1,\dots,\mathfrak{N}\}$ denote the type of the $j$-th jump occurring at time $t_{j}$. The complete measurement record is
$\Gamma_{Q}\equiv\big[(t_{1},m_{1}),(t_{2},m_{2}),\ldots,(t_{J},m_{J})\big],$ (45)
@@
where the subscript Q``Q" is used to emphasize that $\Gamma_{Q}$ is a quantum trajectory. In what follows, variables and functions with subscript $Q$ denote those in the quantum system. The sequence $\Gamma_{Q}$ characterizes one realization of the monitored dynamics and is the basic object from which we construct trajectory-dependent observables. Consider a trajectory observable $N_{Q}(\Gamma_{Q})$, which is a quantum generalization of Eq. (5):
$N_{Q}(\Gamma_{Q})=\sum_{m}C_{m}N_{m},$ (46)
where $N_{m}$ is the number of $m$-th jumps within $0,\tau]$ and $C_{m}$ denotes the corresponding real weight. Let $\Gamma_{Q\varnothing}$ be a quantum trajectory in which there is no jump within the interval $0,\tau]$. Analogously to the condition of Eq. (13), we define the observable $N_{Q\circ}(\Gamma_{Q})$ that satisfies the following condition:
$N_{Q\circ}(\Gamma_{Q\varnothing})=0.$ (47)
$N_{{Q}\circ}$ includes counting observables, which simply count the number of jumps. Let $N_{Q\bullet}(\Gamma_{Q})$ be an arbitrary function defined on $\Gamma_{Q}$. If we denote the sets of functions $N_{Q}$, $N_{Q\circ}$, and $N_{Q\bullet}$ by $\{N_{Q}\}$, $\{N_{Q\circ}\}$, and $\{N_{Q\bullet}\}$, respectively, then inclusion holds:
$\{N_{Q}\}\subset\{N_{Q\circ}\}\subset\{N_{Q\bullet}\}.$ (48)
In a similar manner to the classical case, we consider $K$ replicas of the identical quantum Markov process. Specifically, $\Gamma_{Q}^{(\otimes K)}$ can be constructed from $\left(\Gamma_{Q}^{(1)},\Gamma_{Q}^{(2)},\cdots,\Gamma_{Q}^{(K)}\right)$, where $\Gamma_{Q}^{(k)}$ is a trajectory in the $k$-th process.
II.8 Quantum replica trade-off relations
As in the classical cases, it is straightforward to derive trade-off relations in the quantum replica process. Specifically, it is known that the following relation holds [44]:
@@ II.8 Quantum replica trade-off relations
$\frac{\mathrm{Var}[N_{Q\mathrm{\bullet}}]}{\tau^{2}\left(\partial_{\tau}\mathbb{E}[N_{Q\bullet}]\right)^{2}}\geq\frac{1}{\mathcal{B}(\tau)}.$ (49)
Moreover, when the system is in a steady state, Eq. (49) reduces to [17]
$\frac{\mathrm{Var}[N_{Q}]}{\mathbb{E}[N_{Q}]^{2}}\geq\frac{1}{\mathcal{B}(\tau)}$ (50)
which holds for the counting observable $N_{Q}(\Gamma_{Q})$ defined in Eq. (46). In Eqs. (49) and (50), $\mathcal{B}(\tau)$ denotes the quantum dynamical activity within the interval $0,\tau] [$17, 44, 64]. The classical dynamical activity [Eq. (16)] quantifies the extent of activity of the dynamics. For classical Markov processes, the state change occurs only when there are jumps. However, for the quantum scenario, the state change can occur even without jumps due to the coherent contribution arising from $i[H,\rho]$ term in Eq. (43). The quantum dynamical activity plays a central role in trade-off relations in quantum thermodynamics [17, 44, 65, 64, 66]. For the explicit expression of $\mathcal{B}(\tau)$, see Appendix F.
In the same way as in Eq. (14), from Eq. (49), we derive the following bound for the $K$-replica system:
@@
$\frac{\mathrm{Var}[N_{Q\mathrm{\bullet}}^{(\otimes K)}]}{\tau^{2}\left(\partial_{\tau}\mathbb{E}[N_{Q\bullet}^{(\otimes K)}]\right)^{2}}\geq\frac{1}{K\mathcal{B}(\tau)}.$ (51)
Similarly to the classical counterpart [Eq. (14)], the right-hand side of Eq. (51) depends on the quantum dynamical activity $\mathcal{B}(\tau)$. The left-hand side of Eq. (51) comprises the observable $N_{Q\bullet}(\Gamma_{Q})$, which is an arbitrary function of a quantum trajectory $\Gamma_{Q}$. Moreover, for the observable $N_{Q\circ}$, which requires the condition given in Eq. (47), the following relation holds (see Appendix D for details):
@@
$\frac{\mathbb{E}\left[|N_{Q\circ}^{(\otimes K)}|^{s}\right]^{r/(s-r)}}{\mathbb{E}\left[|N_{Q\circ}^{(\otimes K)}|^{r}\right]^{s/(s-r)}}\geq\left(1-\cos\left[\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{\mathcal{B}(t)}}{t}dt\right]^{2K}\right)^{-1},$ (52)
where
The derivation of Eq. (52) is based on the bound proved in Ref. [45]. For the classical case, the right-hand side depends solely on the initial dynamical activity $\mathfrak{a}(t=0)$. On the other hand, $\mathcal{B}(\tau)$ depends on the dynamics within the interval $0,\tau]$.
II.9 Entropic bound on quantum trajectory observable
We aim to quantify the uncertainty associated with quantum trajectory observables. To achieve this, we analyze the probability distribution of $N_{Q\bullet}(\Gamma_{Q})$. We derive an upper bound for the Tsallis entropy of this distribution, denoted as $P(N_{Q\bullet})$. Following the classical derivation detailed in Appendix A, we obtain
$\left|\frac{\partial}{\partial\tau}H_{q}^{\mathrm{T}}\left[P\left(N_{Q\bullet}\right)\right]\right|\leq\frac{\sqrt{q\mathcal{B}(\tau)}}{2\tau(q-1)},$ (54)
where $q=2,3,\cdots$. Equation (54) holds for arbitrary observable $N_{Q\bullet}(\Gamma_{Q})$ of the quantum trajectory $\Gamma_{Q}$. Again, for $N_{Q\circ}$, which additionally satisfies the condition of Eq. (47), the following relation holds:
$H_{q}^{\mathrm{T}}[P(N_{Q\circ})]\leq\frac{1}{q-1}\sin\left(\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{q\mathcal{B}(t)}}{t}dt\right)^{2},$ (55)
where $q=2,3,\cdots$ and the following condition should be satisfied:
$0\leq\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{q\mathcal{B}(t)}}{t}dt\leq\frac{\pi}{2}.$ (56)
Equations (54) and (55) are quantum generalizations of Eqs. (24) and (25), respectively, and thus their derivations are the same as those of Eqs. (24) and (25). From Eq. (52), we obtain the following relation (see Appendix E for details):
$H_{q}^{\mathrm{T}}[P(N_{Q\circ})]\leq\frac{1-\cos\left[\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{\mathcal{B}(t)}}{t}dt\right]^{2q}}{q-1},$ (57)
where $q=2,3,\cdots$. Equation (57) is a quantum generalization of Eq. (27). Equations (55) and (57) are similar in that the left-hand side is identical and their right-hand sides comprise the quantum dynamical activity. It can be shown that Eq. (57) provides a tighter bound. In the classical case, the left-hand side of Eq. (28) is the Tsallis entropy of the observable $N_{\circ}$ and the right-hand side depends on the dynamical activity $\mathfrak{a}(t=0)$ at time $t=0$. On the other hand, in the quantum case, the right-hand side of Eq. (57) depends on the time integration of the quantum dynamical activity $\mathcal{B}(t)$.
Figure 3: Illustration of the dependence of the Tsallis entropic bound for state distribution [Eq. (35)]. The Tsallis entropic bound shown in Eq. (35) depends on $\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}$, where $B_{\mu_{0}}$ is the initial state, the local escape rate from $B_{\mu_{0}}$. The local escape rate from $B_{\mu_{0}}$ is the sum of transition rates to which the nodes reachable from the initial state with one jump. Suppose the initial state is $B_{1}$, which is colored with dark red. Then, the reachable states are represented by purple nodes. Therefore, the local escape rate becomes $\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}=W_{21}+W_{31}+W_{41}$.
III Numerical simulation
We perform simulations to numerically validate the bounds for the entropy.
First, we perform numerical calculations for the bound for the trajectory observable [Eq. (28)]. We generate a continuous time Markov process, where the transition rate $W_{\mu^{\prime}\mu}$ and the initial probability distribution $p_{\mu}(0)$ are determined randomly. We consider the counting observable $N(\Gamma)$ [Eq. (5)], which simply counts the number of total jumps within the interval $0,\tau]$, where $\tau$ is determined randomly. The range for these random parameters is shown in the caption of Fig. 4. We repeat this procedure to obtain random realizations and calculate the Tsallis entropy and the dynamical activity. For each random realization, we plot $H_{q}^{\mathrm{T}}[P(N_{\circ})]$ versus $\left(1-e^{-q\tau\mathfrak{a}(t=0)}\right)/(q-1)$, which are the left and right-hand sides of Eq. (28), respectively, by a point in Fig. 4. In Figs. 4(a)-(c), we plot the results for (a) $q=1.5$, (b) $q=2$, and (c) $q=3$, where the dashed line denotes the equality case of Eq. (28). In all cases of $q$, the points are above the dashed line, which numerically verifies Eq. (28). The derived relation numerically holds for $q=1.5$, which is non-integer and is a conjectured bound.
Next, we perform numerical calculations for the bounds of the state distribution [Eqs. (31) and (35)]. We generate a random network and consider a continuous time random walk on the network. The transition rate $W_{\mu^{\prime}\mu}$ and the initial node are chosen randomly. We calculate the probability distribution at time $t=\tau$, where $\tau$ is determined randomly, and calculate the Tsallis entropy (Eq. (35)) and its time derivative (Eq. (31)). The range for these random parameters is shown in the caption of Figs. 5 and 6. For each random realization, we plot $\left|\partial_{\tau}H_{q}^{\mathrm{T}}\left[p_{\mu}\left(\tau\mid\mu_{0}\right)\right]\right|$ versus $\sqrt{q\mathcal{A}(\tau)}/(2\tau(q-1))$ (left and right-hand sides of Eq. (31)) in Fig. 5 and $H_{q}^{\mathrm{T}}[p_{\mu}(\tau\mid\mu_{0})]$ versus $(1-e^{-q\tau\sum_{\mu(\mu\neq\mu_{0})}W_{\mu\mu_{0}}})/(q-1)$ (left and right-hand sides of Eq. (35)) in Fig. 6 as points. In Figs. 5 and 6, the dashed lines indicate the equality cases in Eqs. (31) and (35), respectively. In Figs. 5 and 6, we plot the results for (a) $q=1.5$, (b) $q=2$, and (c) $q=3$. For all values of $\alpha$, the points lie above the dashed line, providing numerical verification of Eqs. (31) and (35).
@@ III Numerical simulation
Figure 4: Numerical verification of the Tsallis entropy bound for trajectory observables shown in Eq. (28). The Tsallis entropy $H_{q}^{\mathrm{T}}[P(N_{\circ})]$ versus $\left(1-e^{-q\tau\mathfrak{a}(t=0)}\right)/\left(q-1\right)$ is plotted for (a) $q=1.5$, (b) $q=2$, and (c) $q=3$. Random realizations are shown with points and the solid line indicates the equality case of Eq. (28). In (a)-(c), classical simulations begin by selecting the number of states $D$ between 22 and $5$. Then, the transition rate $W_{\mu^{\prime}\mu}$ is randomly produced, and $\tau$ is chosen within the range of $0.1$ to 1010. The observable $N_{\circ}$ is the number of total jumps within $0,\tau]$. The Tsallis entropy of the observable is estimated by averaging the results of $10^{4}$ simulations.
Figure 5: Numerical verification of the time derivative bound of Tsallis entropy of state distribution shown in Eq. (31). The time derivative of Tsallis entropy $\left|\partial_{\tau}H_{q}^{\mathrm{T}}\left[p_{\mu}(\tau)\right]\right|$ versus $\sqrt{q\mathcal{A}(\tau)}/\left(2\tau(q-1)\right)$ is plotted for (a) $q=1.5$, (b) $q=2$, and (c) $q=3$. Random realizations are shown with points and the dashed line indicates the equality case of Eq. (31). In (a)-(c), the simulations begin by selecting the number of states $D$ between 22 and $20$. Then, the transition rate $W_{\mu^{\prime}\mu}$ is randomly produced, and $\tau$ is chosen within the range of $0.1$ to 1010.
Figure 6: Numerical verification of the Tsallis entropy bound of state distribution shown in Eq. (35). The Tsallis entropy $H_{q}^{\mathrm{T}}[p_{\mu}(\tau\mid\mu_{0})]$ versus $\left(1-e^{-q\tau\sum_{\mu\left(\mu\neq\mu_{0}\right)}W_{\mu\mu_{0}}}\right)/\left(q-1\right)$ is plotted for (a) $q=1.5$, (b) $q=2$, and (c) $q=3$. Random realizations are shown with points and the dashed line indicates the equality case of Eq. (35). In (a)-(c), the simulations begin by selecting the number of states $D$ between 22 and $20$. Then, the transition rate $W_{\mu^{\prime}\mu}$ is randomly produced, and $\tau$ is chosen within the range of $0.1$ to 1010.
IV Conclusion
We have introduced replica Markov processes as a general method to derive thermodynamic trade-off relations for nonlinear functionals of probability distributions. By considering $K$ independent identical copies of a Markov process and applying activity-based trade-off relations to suitably chosen replica observables, nonlinear quantities of the original single process can be bounded by linear expectations in the replica process.
Focusing on entropic measures of uncertainty, we have derived upper bounds on the time derivative and the value of the Tsallis entropy for general trajectory-observable distributions in terms of dynamical activity [Eqs. (24) and (25)], as well as Tsallis bounds for general trajectory observable controlled by the initial activity [Eqs. (27) and (28)]. We also have obtained analogous bounds for network diffusion, yielding an upper bound on the time derivative and the value of the Tsallis entropy of the position distribution [Eqs. (31) and (33)] as well as Tsallis entropy bounds for the state distribution controlled by the initial activity [Eqs. (34) and (35)]. In addition, we illustrated that replicas can treat other nonlinear statistics such as extreme-value observables [Eqs. (39)–(41)]. We extended the construction to continuously monitored open quantum systems, where the bounds are expressed via the quantum dynamical activity [Eqs. (51)–(57)].
These results provide entropic counterparts to activity-based uncertainty relations and establish replica processes as a powerful tool for bounding nonlinear information-theoretic quantities in stochastic and quantum thermodynamics.
Appendix A Derivation of Eqs. (24)–(25) and (31)–(33)
Let us consider the observable:
$\Lambda_{\bullet}^{(\otimes K)}(\Gamma^{(\otimes K)})\equiv\begin{cases}1&N_{\bullet}^{(1)}(\Gamma^{(1)})=N_{\bullet}^{(2)}(\Gamma^{(2)})=\cdots=N_{\bullet}^{(K)}(\Gamma^{(K)})\\ 0&\mathrm{otherwise}\end{cases}.$ (58)
$\Lambda_{\bullet}^{(\otimes K)}$ returns 1 only when the values of the observable $N_{\bullet}$ in the $K$ processes agree, $N_{\bullet}^{(1)}(\Gamma^{(1)})=N_{\bullet}^{(2)}(\Gamma^{(2)})=\cdots=N_{\bullet}^{(K)}(\Gamma^{(K)})$. Let $\mathcal{Z}_{\bullet}$ be the set of all possible values that the observable $N_{\bullet}(\Gamma)$ can take. The expectation and variance of Eq. (58) are
$\mathbb{E}[\Lambda_{\bullet}^{(\otimes K)}(\Gamma^{\otimes K})]$ $=\sum_{z\in\mathcal{Z}_{\bullet}}P_{\bullet}(z)^{K},$ (59)
$\mathrm{Var}[\Lambda_{\bullet}^{(\otimes K)}(\Gamma^{\otimes K})]$ $=\sum_{z\in\mathcal{Z}_{\bullet}}P_{\bullet}(z)^{K}\left(1-\sum_{z\in\mathcal{Z}_{\bullet}}P_{\bullet}(z)^{K}\right).$ (60)
where we abbreviated $P_{\bullet}(z)\equiv P(N_{\bullet}=z)$. Substituting Eqs. (59) and (60) into Eq. (14), we obtain
$\frac{\left|\partial_{\tau}\sum_{z}P_{\bullet}(z)^{K}\right|}{\sqrt{\sum_{z}P_{\bullet}(z)^{K}\left(1-\sum_{z}P_{\bullet}(z)^{K}\right)}}\leq\frac{\sqrt{K\mathcal{A}(\tau)}}{\tau}.$ (61)
For an arbitrary function $f(\tau)$, the following relation holds:
$\frac{d}{d\tau}\arcsin\left(2f-1\right)=\frac{1}{\sqrt{f(1-f)}}\left(\frac{df}{d\tau}\right).$ (62)
Using Eq. (62) in Eq. (61), we have
$\left|\frac{d}{d\tau}\arcsin\left(2\sum_{z}P_{\bullet}(z)^{K}-1\right)\right|\leq\frac{\sqrt{K\mathcal{A}(\tau)}}{\tau}.$ (63)
Recall that the following relation holds for $|x|<1$:
$\left|\frac{d}{dx}\left(x\right)\right|\leq\left|\frac{d}{dx}\arcsin\left(x\right)\right|.$ (64)
Using Eq. (64) with Eq. (63) yields
$\left|\frac{d}{d\tau}\sum_{z}P_{\bullet}(z)^{K}\right|\leq\frac{\sqrt{K\mathcal{A}(\tau)}}{2\tau},$ (65)
which proves Eq. (24) in the main text.
Next, we derive Eq. (25). Instead of $N_{\bullet}$, we consider $N_{\circ}$ in Eq. (63). While $N_{\bullet}$ is an arbitrary function of $\Gamma$, $N_{\circ}$ has the additional assumption given in Eq. (6). Let $\mathcal{Z}_{\circ}$ be the set of all possible values that the observable $N_{\circ}(\Gamma)$ can take. By integrating Eq. (63) from $t=0$ to $t=\tau$ and using the triangle inequality, we obtain
$\left|\arcsin\left(2\sum_{z\in\mathcal{Z}_{\circ}}P(N_{\circ}=z)^{K}-1\right)-\frac{\pi}{2}\right|$
$\leq\int_{0}^{\tau}\frac{\sqrt{K\mathcal{A}(t)}}{t}dt.$ (66)
In Eq. (66), we used the fact that $N_{\circ}=0$ at $t=0$ due to the condition of Eq. (6) (recall that, at time $t=0$, there is no jump). Therefore, the probability is concentrated on $P(N_{\circ}=0)=1$ at time $t=0$. When $0\leq(1/2)\int_{0}^{\tau}\sqrt{K\mathcal{A}(t)}/t\,dt\leq\pi/2$, we obtain from Eq. (66)
$\sum_{z\in\mathcal{Z}_{\circ}}P(N_{\circ}=z)^{K}\geq\cos\left(\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{K\mathcal{A}(t)}}{t}dt\right)^{2},$ (67)
which yields Eq. (25) in the main text.
By adopting the derivation above, we can obtain the bounds shown in Eqs. (31)–(33). Since $N_{\bullet}(\Gamma)$ can be an arbitrary function of $\Gamma$, we can take
$N_{\bullet}(\Gamma)=X(\tau),$ (68)
where $X(\tau)$ is the state of the Markov process at time $t=\tau$. Using Eq. (68) in Eq. (65), we obtain
$\left|\frac{d}{d\tau}\sum_{\mu}p_{\mu}(\tau)^{K}\right|\leq\frac{\sqrt{K\mathcal{A}(\tau)}}{2\tau},$ (69)
Moreover, integrating Eq. (69) from $t=0$ to $t=\tau$, we obtain
$\left|\sum_{\mu}p_{\mu}(\tau)^{K}-\sum_{\mu}p_{\mu}(0)^{K}\right|\leq\int_{0}^{\tau}\frac{\sqrt{K\mathcal{A}(t)}}{2t}dt,$ (70)
which derives Eq. (32) in the main text. When the initial state of the process is fixed to $B_{\mu_{0}}$, indicating $p_{\mu_{0}}(0)=1$, we can derive a tighter inequality than Eq. (70). Specifically, substituting Eq. (68) into Eq. (63), we obtain
$\left|\arcsin\left(2\sum_{\mu}p_{\mu}(\tau\mid\mu_{0})^{K}-1\right)-\frac{\pi}{2}\right|\leq\int_{0}^{\tau}\frac{\sqrt{K\mathcal{A}(t)}}{t}dt,$ (71)
which derives
$\sum_{\mu}p_{\mu}(\tau\mid\mu_{0})^{K}\geq\cos\left(\frac{1}{2}\int_{0}^{\tau}\frac{\sqrt{K\mathcal{A}(t)}}{t}dt\right)^{2},$ (72)
where $0\leq(1/2)\int_{0}^{\tau}\sqrt{K\mathcal{A}(t)}/t\,dt\leq\pi/2$ should be satisfied. Equation (72) derives Eq. (33) in the main text.
Appendix B Derivation of Eq. (17)